A plant virus is found to consist of uniform cylindrical particle for 150 A in diameter A long . The specific volume of virus is 0.75 cm3/gm . If the virus is considered to be a single particle find its molarity

Grade:11

1 Answers

Tony
5 years ago
Diameter of virus (d) = 150 oA
​Length of virus =5000 oA

Specific volume  =0.75 cm3/g

Volume of a virus = πr2l =  3.14 x [(150/2) x 10-8]2 x 5000 x 10-8 = 8.8312 x 10-17 cm3

Mass of 1 virus = 
0.884 *10-16/0.75 g
= 1.178 * 10-19 g
∴ Mol. wt. of virus = = 1.178 * 10-19 * 6.02 * 1023
= 7.09 * 107
8.8312 x 10−170.75=1.1775



1.1775 ×1016 ×6.023 × 1023 =7.0920 × 107